OpenStudy (anonymous):

Which of the following are solutions to the equation below? 9x2 + 9x + 1 = 19

OpenStudy (anonymous):

x=-2,x=1

OpenStudy (anonymous):

9x2+9x+1=19 -19=-19 9x2+9x-18=0 (3x+6)(3x-3) 3x+6=0 | 3x-3=0 -6=-6| +3=+3 3x=-6 | 3x=3 /3 /3 | /3 /3 x=-2 | x=1